OpenStudy (anonymous):

-ax+3b>5 solve for x PLEASE HELP ME

OpenStudy (imstuck):

The best way to do these is to pretend that the inequality sign is an equals sign; takes off some of the pressure. WE can change it back when we are done with it.

OpenStudy (imstuck):

In an equation, you would begin by subtracting the 3b over to the other side, so let's do that:

OpenStudy (imstuck):

\[-ax=5-3b\]

OpenStudy (imstuck):

Now to solve for x you will divide by -a:\[x=\frac{ 5-3b }{ -a }\]

OpenStudy (imstuck):

Now a couple of things here going back to the inequality thing. When you divide by a negative number you know that the direction of the sign changes. So we will do that now:

OpenStudy (imstuck):

\[x <\frac{ 5-3b }{ -a }\]

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

And then the negative in with the a changes the sign in the "5-3b" to a positive one, like this:

OpenStudy (imstuck):

\[x <-\frac{ 5+3b }{ a }\]

OpenStudy (imstuck):

And that's it!

OpenStudy (anonymous):

thank you so much!!

OpenStudy (imstuck):

Youre very welcome!

OpenStudy (anonymous):

can I ask you another question?

OpenStudy (anonymous):

@IMSsuck